# Local Dimension Spectrum for Dominated Irreducible Planar Self-Affine Sets

### v1

@@ Abstract
The local dimension spectrum gives a way to quantify fractal properties of a given measure. It is well understood for self-similar measures without overlaps. In this article we study the local dimension spectrum for dominated self-affine measures. Our main results completely characterize the behavior of the symbolic spectrum, and after detailed analysis on the exact dimensionality of self-affine measures we exhibit new deterministic results for the local dimension spectrum on a much wider range than was previously understood.
@@ 1 Introduction
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a tuple of contractive invertible $2\times 2$-matrices. If $(v_{1},\ldots,v_{N})\in(\mathbb{R}^{2})^{N}$ is a tuple of translation vectors, then the collection $(\varphi_{1},\ldots,\varphi_{N})$ of contractive invertible affine maps given by
@@
The self-affine set $X$ can also be expressed as the image $\pi(\Sigma)$ of the shift space $\Sigma=\{1,\ldots,N\}^{\mathbb{N}}$, where $\pi\colon\Sigma\to X$ is the canonical projection of the IFS and shift refers to the left shift $\sigma\colon\Sigma\to\Sigma$; see Sections 5 and 2.1 respectively for definitions. If $\phi\colon\Sigma\to\mathbb{R}$ is a continuous function and $\mu$ is a Borel probability measure, ergodic with respect to $\sigma$, then by the Birkhoff ergodic theorem,
@@
for $\mu$-almost all $\mathtt{i}\in\Sigma$. The Birkhoff ergodic theorem does not give any information about the exceptional set where this equality does not hold. Since any two ergodic measures are mutually singular the exceptional set contains other ergodic measures and thus, the exceptional set cannot be considered negligible. Bárány, Jordan, Käenmäki, and Rams [4, Theorem 2.1] managed to understand the structure of this exceptional set on self-affine sets. They determined the Hausdorff dimension of the canonical projection of the set of $\mathtt{i}\in\Sigma$ for which the Birkhoff averages $\frac{1}{n}\sum_{j=0}^{n-1}\phi(\sigma^{j}\mathtt{i})$ converge to a given value. As a consequence, they were also able to determine the Hausdorff dimension spectrum of the Lyapunov exponents. A much harder problem on self-affine sets is to understand the local dimension spectrum.
@@
provided the limit exists. If the limit does not exist, then the corresponding upper and lower limits are denoted by $\operatorname{\overline{dim}_{loc}}(\mu,x)$ and $\operatorname{\underline{dim}_{loc}}(\mu,x)$, respectively. The local dimension of $\mu$ is intrinsically connected to the dimension of the subsets of $X$: one can see rather directly that
@@
We are interested in determining the Hausdorff dimension of these level sets. In other words, we wish to determine the local dimension spectrum of $\mu$ by which we mean the behavior of the function $s\mapsto\operatorname{dim_{H}}(X(\mu,s))$. The local dimension spectrum gives an important way of quantifying fractal properties of the measure $\mu$. It is well understood for self-similar measures without overlaps and some progress has been made in understanding both the overlapping self-similar and the self-affine cases.
In particular, given a tuple of matrices $(A_{1},\ldots,A_{N})$, each with $\|A_{i}\|<\frac{1}{2}$, and a probability vector $(p_{1},\ldots,p_{N})$, Barral and Feng [8], building on work of Falconer [13], were able to give a formula for part of the corresponding dimension spectrum which holds for almost every set of translation vectors $v_{1},\ldots,v_{N}$. Given $q\neq 1$, we let $s_{q}$ be the unique zero of an associated pressure function $s\mapsto P(\psi^{q,s})$ defined in Section 3. The $L^{q}$-spectrum $\tau$ is defined and its properties are studied in Section 4. As $\tau$ is concave, we denote its left and right derivatives by $\tau_{\pm}^{\prime}(q)$ with the convention that if in one occurrence it means the left (or respectively right) derivative, then it means that everywhere.
Theorem 1.1 (Barral-Feng [8, Theorem 1.3]).
Suppose that $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is such that $\|A_{i}\|<\frac{1}{2}$ for all $i\in\{1,\ldots,N\}$ and $(p_{1},\ldots,p_{N})$ is a probability vector. Given a choice of translation vector $\mathbf{v}=(v_{1},\ldots,v_{N})$, let $\mu_{\mathbf{v}}$ be the corresponding self-affine measure. If $10<q<$ 1, $s_{q}<1$, and $q\tau_{\pm}^{\prime}(q)-\tau(q)\leqslant 1$, then
@@
This almost everywhere result uses the transversality technique and is very much in the spirit of earlier work of Falconer giving an almost everywhere result for the Hausdorff dimension of $\mu$; see [12]. Barral and Feng also have results for $q>1$ in the case that the matrices $A_{i}$ are all diagonal.
In recent years there has been very substantial progress in understanding the Hausdorff dimension of self-affine measures. In the 1980s Falconer [12] gave a formula for the expected dimension of a self-affine measure and showed that this formula does indeed give the dimension typically, but there were rather few explicit examples of self-affine measures satisfying Falconer’s formula until 2010s. More recently, sufficient conditions for Falconer’s formula to be correct were given in terms of properties of orthogonal projections of the self-affine measure [2, 5, 14], and these in turn were used to show that wide classes of self-affine measures satisfy Falconer’s formula [3].
Our goal in this article is to mirror some of the recent progress on Hausdorff dimension, and to replace the “almost-everywhere” in results of [8] quoted above with explicit projection results, as well as extending the range of applicability of those results. We also consider much larger class of quasi-Bernoulli measures rather than just self-affine measures, which are canonical projections of Bernoulli measures, and in Theorem 4.3 we completely characterize the behavior of the symbolic spectrum.
A planar self-affine set $X$ is dominated and irreducible if the linear parts of the defining affine maps of $X$ have positive entries and they do not share a common invariant line; see Sections 2.3 and 5 respectively for precise definitions. The set of Furstenberg directions $X_{F}$ is the closure or all top Oseledets directions $V$ for the inverse linear action; see Section 2.3 for a precise definition. Finally, we say that $X$ satisfies the projective strong separation condition if
@@
for all $V\in X_{F}$ whenever $i\neq j$. It is evident that the projective strong separation condition implies the projective open set condition defined in [7, §3]. It also implies the strong separation condition which means that $\varphi_{i}(X)\cap\varphi_{j}(X)=\emptyset$ whenever $i\neq j$.
Our first main result covers the dimension range $(0,1)$.
@@
Suppose that $X$ is a dominated irreducible planar self-affine set satisfying the projective strong separation condition and $\mu$ is a fully supported quasi-Bernoulli measure on $X$. If $0<q\neq 1$, $10<q\tau^{\prime}(q)-\tau(q)<$ 1, and $10<\tau(q)/(q-1)<$ 1, then
@@
The second main result covers the dimension range $(1,2)$.
@@
Suppose that $X$ is a dominated irreducible planar self-affine set satisfying the strong separation condition and $\mu$ is a fully supported quasi-Bernoulli measure on $X$ such that $(\operatorname{proj}_{V^{\bot}})_{*}\mu$ is absolutely continuous with uniformly bounded density over all $V\in X_{F}$. If $0<q\neq 1$, $21<q\tau^{\prime}(q)-\tau(q)<$ 2, and $21<\tau(q)/(q-1)<$ 2, then
@@
These theorems are proved via a sequence of propositions stated in Section 6. The propositions in Section 6 are slightly more general than Theorems 1.2 and 1.3, for example the proof of the lower bound of Theorem 1.3 requires only that the projections of $\mu$ have dimension one, not that they are absolutely continuous with bounded density.
The remainder of the paper is organized as follows: In Sections 24 we work in the shift space and cover necessary preliminaries on thermodynamic formalism and determine the properties of the symbolic $L^{q}$-spectrum. In Sections 5 and 6 we work in the plane and study exact dimensionality of quasi-Bernoulli measures in more detail and transfer the symbolic result into self-affine sets.
@@ 2.1 Shift Space
Let $\Sigma=\{1,\ldots,N\}^{\mathbb{N}}$ be the collection of all infinite words obtained from alphabet $\{1,\ldots,N\}$. If $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$, then we define $\mathtt{i}|_{n}=i_{1}\cdots i_{n}$ for all $n\in\mathbb{N}$. If $\mathtt{i}=i_{1}\cdots i_{n}$, then we write $\overleftarrow{\mathtt{i}}=i_{n}\cdots i_{1}$. The empty word $\mathtt{i}|_{0}$ is denoted by $\varnothing$. Define $\Sigma_{n}=\{\mathtt{i}|_{n}:\mathtt{i}\in\Sigma\}$ for all $n\in\mathbb{N}$ and $\Sigma_{*}=\bigcup_{n\in\mathbb{N}}\Sigma_{n}\cup\{\varnothing\}$. Thus $\Sigma_{*}$ is the collection of all finite words. The length of $\mathtt{i}\in\Sigma_{*}\cup\Sigma$ is denoted by $|\mathtt{i}|$. The longest common prefix of $\mathtt{i},\mathtt{j}\in\Sigma_{*}\cup\Sigma$ is denoted by $\mathtt{i}\wedge\mathtt{j}$ and the concatenation of two words $\mathtt{i}\in\Sigma_{*}$ and $\mathtt{j}\in\Sigma_{*}\cup\Sigma$ is denoted by $\mathtt{i}\mathtt{j}$. Let $\sigma$ be the left shift operator defined by $\sigma\mathtt{i}=i_{2}i_{3}\cdots$ for all $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$. If $\mathtt{i}\in\Sigma_{n}$ for some $n$, then we set $[\mathtt{i}]=\{\mathtt{j}\in\Sigma:\mathtt{j}|_{n}=\mathtt{i}\}$. The set $[\mathtt{i}]$ is called a cylinder set.
If $\mu\in\mathcal{M}(\Sigma)$, where $\mathcal{M}(\Sigma)$ denote the collection of all Borel probability measures on $\Sigma$, then the pushforward measure of $\nu$ under a measurable map $f$ defined on $\Sigma$ is denoted by $f_{*}\nu$. We say that a measure $\mu\in\mathcal{M}(\Sigma)$ is fully supported if each cylinder set has positive measure.

### v2

# Local Dimension Spectrum for Dominated Planar Self-Affine Sets

@@ Abstract
The local dimension spectrum provides a framework for quantifying the fractal properties of a measure, and it is well understood for non-overlapping self-similar measures. In this article, we study the local dimension spectrum for dominated self-affine measures. By analyzing exact dimensionality, we obtain deterministic results that extend the scope of the local dimension spectrum beyond the almost-sure setting.
@@ 1 Introduction
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a tuple of contractive invertible $2\times 2$-matrices. Given a tuple of translation vectors $(v_{1},\ldots,v_{N})\in(\mathbb{R}^{2})^{N}$, the collection $(\varphi_{1},\ldots,\varphi_{N})$ of contractive invertible affine maps defined by
@@
The self-affine set $X$ can be expressed as the image $\pi(\Sigma)$ of the shift space $\Sigma=\{1,\ldots,N\}^{\mathbb{N}}$, where $\pi\colon\Sigma\to X$ is the canonical projection of the IFS and $\sigma\colon\Sigma\to\Sigma$ is the left shift; see Sections 5 and 2.1 for the corresponding definitions. If $\phi\colon\Sigma\to\mathbb{R}$ is a continuous function and $\mu$ is a Borel probability measure ergodic with respect to $\sigma$, then by the Birkhoff ergodic theorem,
@@
for $\mu$-almost all $\mathtt{i}\in\Sigma$. The Birkhoff ergodic theorem does not provide information about the exceptional set on which this equality fails. Since any two distinct ergodic measures are mutually singular, the exceptional set supports other ergodic measures and therefore cannot be regarded as negligible. Bárány, Jordan, Käenmäki, and Rams [3, Theorem 2.1] analyzed the structure of this exceptional set for self-affine sets. They determined the Hausdorff dimension of the canonical projection of the set of $\mathtt{i}\in\Sigma$ for which the Birkhoff averages $\frac{1}{n}\sum_{j=0}^{n-1}\phi(\sigma^{j}\mathtt{i})$ converge to a given value. As a consequence, they were also able to determine the Hausdorff dimension spectrum of the Lyapunov exponents. A much harder problem on self-affine sets is to understand the local dimension spectrum.
@@
provided the limit exists. If the limit does not exist, the corresponding upper and lower limits are denoted by $\operatorname{\overline{dim}_{loc}}(\mu,x)$ and $\operatorname{\underline{dim}_{loc}}(\mu,x)$, respectively. The local dimension of $\mu$ is intrinsically connected to the dimension of subsets of $X$. It is easy to see that
@@
We are interested in determining the Hausdorff dimension of these level sets. The $L^{q}$-spectrum of $\mu$ is defined as
$\tau(\mu,q)=\liminf_{r\downarrow 0}\frac{\log\sup\sum_{i}\mu(B(x_{i},r))^{q}}{\log r},$
where the supremum is taken over all collections of pairwise disjoint balls $\{B(x_{i},r)\}_{i}$ centered in the support of $\mu$. We say that the multifractal formalism holds for $\mu$ at $s$ if
$\operatorname{dim_{H}}(X(\mu,s))=\inf\{qs-\tau(\mu,q):q\in\mathbb{R}\}.$
Our goal is to investigate the local dimension spectrum of $\mu$, that is, the behavior of the function $s\mapsto\operatorname{dim_{H}}(X(\mu,s))$, and to provide tools to assess when the multifractal formalism holds for $\mu$. This spectrum is well understood for self-similar measures satisfying the open set condition, with some progress in the overlapping self-similar and self-affine cases. Notable results on self-similar sets and self-affine carpets involving exact overlaps include [25, 30, 11, 21, 9]; results for Hueter-Lalley type self-affine sets are given in [20].
In particular, for a tuple of matrices $(A_{1},\ldots,A_{N})$, each with $\|A_{i}\|<\frac{1}{2}$, and a probability vector $(p_{1},\ldots,p_{N})$, Barral and Feng [10], building on Falconer’s work [18], derived a formula for part of the corresponding local dimension spectrum, valid for almost every choice of translation vectors $v_{1},\ldots,v_{N}$. Falconer [18] established (see [10, Theorem 1.2]) that the symbolic $L^{q}$-spectrum $\tau$, defined in Section 4, coincides with the $L^{q}$-spectrum for almost all translation vectors when $1<q\leqslant 2$. Since $\tau$ is concave, we write $\tau_{\pm}^{\prime}(q)$ for its left and right derivatives. We adopt the convention that whenever a one-sided derivative is used at some $q$, the same choice is used throughout.
Theorem 1.1 (Barral-Feng [10, Theorem 1.3]).
Suppose that $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is such that $\|A_{i}\|<\frac{1}{2}$ for all $i\in\{1,\ldots,N\}$ and $(p_{1},\ldots,p_{N})$ is a probability vector. Given a choice of translation vector $\mathbf{v}=(v_{1},\ldots,v_{N})$, let $\mu_{\mathbf{v}}$ be the corresponding self-affine measure. If $0<q<1$, $0<q\tau_{\pm}^{\prime}(q)-\tau(q)<1$, and $0<\tau(q)/(q-1)<1$, then $\tau(\mu_{\mathbf{v}},q)=\tau(q)$ and
@@
This almost-sure result relies on a transversality argument, in the spirit of Falconer’s work on the almost-sure Hausdorff dimension of self-affine sets; see [17]. For $q>1$, Barral and Feng provide analogous results when the matrices $A_{i}$ are diagonal. According to [10, §7], the results of Barral and Feng extend to a broader class of quasi-Bernoulli measures, beyond self-affine measures (which are canonical projections of Bernoulli measures).
In recent years there has been substantial progress in understanding the Hausdorff dimension of self-affine sets and measures. In the 1980s Falconer [17] gave a formula for the generic dimension of a self-affine set and showed that this formula holds typically. However, there were few explicit examples of self-affine sets satisfying Falconer’s formula until the 2010s. More recently, sufficient conditions for the validity of Falconer’s formula were given in terms of properties of orthogonal projections of the self-affine measure; see [7, 6, 16, 19]. These were then used to show that wide classes of self-affine sets satisfy Falconer’s formula; see [2, 32, 28].
Our goal is to mirror recent progress on the Hausdorff dimension, replacing the “almost-everywhere” statements in [10] with explicit projection results and extending their range of applicability. The following proposition completely characterizes the behavior of the symbolic spectrum. Here $\operatorname{dim_{L}}(\eta)$ is the Lyapunov dimension of a shift-invariant measure $\eta$ and $\operatorname{dim_{L}}(\mu,\eta)$ is the Lyapunov cross-dimension of $\mu$ relative to $\eta$; see Section 2.2. A more detailed statement can be found in Theorem 4.3.
Proposition 1.2.
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices, $\mu$ be a fully supported quasi-Bernoulli measure on $\Sigma$, and $q\neq 1$. If $\tau^{\prime}(q)$, $\tau(q)/(q-1)$, and $q\tau^{\prime}(q)-\tau(q)$ are contained in the same interval, $(0,1)$, $(1,2)$, or $(2,\infty)$, then $\tau^{\prime}(q)\leqslant\tau(q)/(q-1)$ and
$\sup\{\operatorname{dim_{L}}(\eta):\eta\text{ is such that }\operatorname{dim_{L}}(\mu,\eta)=\tau^{\prime}(q)\}=q\tau^{\prime}(q)-\tau(q).$
Note that the interval containing $\tau(q)/(q-1)$ determines which singular values must be considered. The Lyapunov dimension of the auxiliary measure that attains the above supremum and the Lyapunov cross-dimension of $\mu$ relative to $\eta$ are identified via $q\tau^{\prime}(q)-\tau(q)$ and $\tau^{\prime}(q)$, respectively, in the same interval as $\tau(q)/(q-1)$.
The Furstenberg direction set $X_{F}$ is the closure of the top Oseledets directions $V$ for the inverse linear action; see Section 2.3 for details. A planar self-affine set $X$ is dominated, for instance, when the linear parts of its defining affine maps have positive entries; see Section 2.3 for a precise definition. Throughout this work, we assume that $X$ is dominated and that $X_{F}$ is not a singleton. This holds, for example, when a dominated self-affine set is irreducible, meaning that the linear parts of its defining affine maps share no common invariant line; see Section 5. Finally, we say that $X$ satisfies the projective strong separation condition if
@@
for all $V\in X_{F}$ whenever $i\neq j$, where $\operatorname{proj}_{V^{\perp}}$ is the orthogonal projection onto $V^{\perp}$. It is evident that the projective strong separation condition implies the projective open set condition defined in [5, §3]. It also implies the strong separation condition, meaning that $\varphi_{i}(X)\cap\varphi_{j}(X)=\emptyset$ whenever $i\neq j$. Our first main result covers the dimension range $(0,1)$.
@@
Suppose that $X$ is a dominated planar self-affine set satisfying the projective strong separation condition such that $X_{F}$ is not a singleton, and that $\mu$ is a fully supported quasi-Bernoulli measure on $X$. If $0<q\neq 1$, $0<q\tau^{\prime}(q)-\tau(q)<1$, and $0<\tau(q)/(q-1)<1$, then
@@
To gain further insight into the projective strong separation condition, observe that $\operatorname{dim_{H}}(X)<1$ since $X_{F}$ is invariant. Hence, this condition implies all equivalent statements in [5, Theorem 3.5]. In particular, it ensures that $\operatorname{proj}_{V^{\bot}}(X)$ is Ahlfors regular for all $V\in X_{F}$; thus, measures supported on $\operatorname{proj}_{V^{\bot}}(X)$ cannot be absolutely continuous. Examples of dominated, irreducible planar self-affine sets satisfying the projective strong separation condition are given in [5, Corollary 1.4]. The second main result covers the dimension range $(1,2)$.
@@
Suppose that $X$ is a dominated planar self-affine set satisfying the strong separation condition such that $X_{F}$ is not a singleton, and that $\mu$ is a fully supported quasi-Bernoulli measure on $X$ such that $(\operatorname{proj}_{V^{\bot}})_{*}\mu$ is absolutely continuous with uniformly bounded density over all $V\in X_{F}$. If $0<q\neq 1$, $1<q\tau^{\prime}(q)-\tau(q)<2$, and $1<\tau(q)/(q-1)<2$, then
@@
Theorems 1.3 and 1.4 are proved via a sequence of propositions stated in Section 6. The propositions in Section 6 are slightly more general than Theorems 1.3 and 1.4; for example, the proof of the lower bound in Theorem 1.4 requires only that the projections of $\mu$ have dimension one, not that they are absolutely continuous with bounded density.
Recall that, by [5, Lemma 2.10] and [2, Theorem 7.1], we have
$\operatorname{\underline{dim}_{H}}((\operatorname{proj}_{V^{\perp}})_{*}\mu)=1$
for all $V\in\mathbb{RP}^{1}$ and fully supported self-affine measures $\mu$ on $X$ with $\operatorname{\underline{dim}_{H}}(\mu)\geqslant 1$. Recently, Bárány [8, Theorem 1.2(i)] strengthened this result for the equilibrium state of the singular value function. Specifically, if $\mathcal{H}^{s}(X)>0$, where $1<s\leqslant 2$ is the affinity dimension, then the equilibrium state satisfies the conditions of Theorem 1.4. The affinity dimension and the equilibrium state for the singular value function are defined in Section 2.4. The following proposition extends this by adapting the techniques of [8] to construct a class of self-affine measures satisfying the conditions of Theorem 1.4.
Proposition 1.5.
Let $\varphi_{i}\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$, $\varphi_{i}(x)=A_{i}x+v_{i}$, where $v_{i}=(v_{i}^{1},v_{i}^{2})$ and
$A_{i}=\begin{pmatrix}a_{i}&0\\ b_{i}&c_{i}\end{pmatrix}$
with $0<|a_{i}|<|c_{i}|<\tfrac{1}{2}$ for all $i\in\{1,\ldots,N\}$ such that the matrices $A_{i}$ are not simultaneously diagonalizable. If the associated self-affine set $X$ satisfies the strong separation condition and $\mu$ is a fully supported quasi-Bernoulli measure satisfying
$\inf\biggl\{s\geqslant 0:\sum_{n\in\mathbb{N}}\sum_{\mathtt{i}\in\Sigma_{n}}\frac{\mu([\mathtt{i}])^{2}}{\|A_{\mathtt{i}}\|^{s}}<\infty\biggr\}>3,$ (1.3)
then $\mathsf{A}=(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})$ is dominated, $X_{F}$ is not a singleton, and $(\operatorname{proj}_{V^{\bot}})_{*}\mu$ is absolutely continuous with uniformly bounded density over all $V\in X_{F}$ for Lebesgue almost every choice of $(v_{1}^{2},\ldots,v_{N}^{2})$.
We prove Proposition 1.5 at the end of Section 6. This proposition confirms that the class of self-affine sets and measures satisfying the conditions of Theorem 1.4 is non-empty. We also remark that verifying condition (1.3) for self-affine measures is straightforward; see (6.12). Given that this result relies on a random choice of translation vectors, constructing explicit examples remains a challenging and open problem.
The paper is organized as follows. In Sections 2 and 4, we work in the shift space, cover preliminaries on thermodynamic formalism, and determine the properties of the symbolic $L^{q}$-spectrum. In Sections 5 and 6, we work in the plane, study the exact dimensionality of quasi-Bernoulli measures in detail, and transfer the symbolic results to self-affine sets.
@@ 2.1 Shift Space
Let $\Sigma=\{1,\ldots,N\}^{\mathbb{N}}$ be the collection of all infinite words obtained from the alphabet $\{1,\ldots,N\}$. If $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$, then we define $\mathtt{i}|_{n}=i_{1}\cdots i_{n}$ for all $n\in\mathbb{N}$. If $\mathtt{i}=i_{1}\cdots i_{n}$, then we write $\overleftarrow{\mathtt{i}}=i_{n}\cdots i_{1}$. The empty word $\mathtt{i}|_{0}$ is denoted by $\varnothing$. Define $\Sigma_{n}=\{\mathtt{i}|_{n}:\mathtt{i}\in\Sigma\}$ for all $n\in\mathbb{N}$ and $\Sigma_{*}=\bigcup_{n\in\mathbb{N}}\Sigma_{n}\cup\{\varnothing\}$. Thus $\Sigma_{*}$ is the collection of all finite words. The length of $\mathtt{i}\in\Sigma_{*}\cup\Sigma$ is denoted by $|\mathtt{i}|$. The longest common prefix of $\mathtt{i},\mathtt{j}\in\Sigma_{*}\cup\Sigma$ is denoted by $\mathtt{i}\wedge\mathtt{j}$ and the concatenation of two words $\mathtt{i}\in\Sigma_{*}$ and $\mathtt{j}\in\Sigma_{*}\cup\Sigma$ is denoted by $\mathtt{i}\mathtt{j}$. Let $\sigma$ be the left shift operator defined by $\sigma\mathtt{i}=i_{2}i_{3}\cdots$ for all $\mathtt{i}=i_{1}i_{2}\cdots\in\Sigma$. If $\mathtt{i}\in\Sigma_{n}$ for some $n$, then we set $\mathtt{i}]=\{\mathtt{j}\in\Sigma:\mathtt{j}|_{n}=\mathtt{i}\}$. The set $\mathtt{i}]$ is called a cylinder set.
If $\mu\in\mathcal{M}(\Sigma)$, where $\mathcal{M}(\Sigma)$ denotes the collection of all Borel probability measures on $\Sigma$, then for a measurable map $f$ defined on $\Sigma$ we denote the pushforward of a measure $\nu$ by $f_{*}\nu$. We say that a measure $\mu\in\mathcal{M}(\Sigma)$ is fully supported if every cylinder set has positive measure.